How do I find the x and the y, using elimination, when given the equations -4x-15y=-17, and -x+5y=-13?
Answer by Boreal(15235)

-4x-15y=-17
-x+5y=-13
multiply the second by 3
-3x+15y=-39
now add the two, and the y will disappear
-7x=-56
x=8
y=-1
(8, -1)
check in both